alones (alones):

@satellite73 @satellite73 @satellite73 I GOT TO MIDPOINT FINALLY @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73

satellite73 (satellite73):

what !!!

alones (alones):

i think i forgot how to olve the midpoint, tho i stil remember the formula /.< Find the midpoint of the line segment whose endpoints are (-2, 5) and (4, -9). i keep on getting fraction >.>

satellite73 (satellite73):

\[\huge \left(\frac{x_1+x_2}{2},\frac{y_1+y_2}{2}\right)\]

alones (alones):

yes this.. i alreadys tudied the formula.. and still keep on stucking

satellite73 (satellite73):

\[(-2, 5) , (4, -9)\]\[\left(\frac{4-2}{2},\frac{5-9}{2}\right)\]

alones (alones):

this is how i wrote\[\frac{ -2+4 }{ 2 } \frac{ 5+-9 }{ 2 }\]

satellite73 (satellite73):

ok same difference as the kids say

alones (alones):

i added them, wait so i suppose to subtract?

satellite73 (satellite73):

no do what you wrote exactly

alones (alones):

waht i wrote exactly?

satellite73 (satellite73):

yes just compute it

alones (alones):

i got (1,7)

satellite73 (satellite73):

the one is right

satellite73 (satellite73):

\[\frac{5+(-9)}{2}=?\]

alones (alones):

damn, i don't have it as

alones (alones):

-2?

satellite73 (satellite73):

yes

alones (alones):

wait so what? (1,-2) leik this?

satellite73 (satellite73):

eys like that

alones (alones):

jeez satellite.. sorry my key is formed wrongly.. it will take me a log time to re-arreange them.. so basically i type as i'm used to on a normal one @satellite73 ayee

alones (alones):

adn thansk

satellite73 (satellite73):

??

alones (alones):

Oh nothing :X

alones (alones):

@CandyCove LOL WHY DID U DELETE THAT x'D